OpenStudy (anonymous):

A high-altitude spherical weather balloon expands as it rises due to the drop in atmospheric pressure. Suppose that the radius r increases at the rate of 0.16 inches per second and that r = 38 inches at time t = 0. Determine the equation that models the volume V of the balloon at time t and find the volume when t = 280 seconds.

OpenStudy (anonymous):

V(t) = 4pi(38+0.16t)^3/3 ; 1,129,910.14 in.3 V(t) = 4pi(38+0.16t)^3 ; 2,377,823.53 in.3 V(t) = 4π(0.16t)2; 25,221.21 in.3 V(t) = 4π(38 + 0.16t)2; 1,325,689.77 in.3

OpenStudy (ranga):

V = 4/3*pi*r^3 Take the derivative with respect to t dV/dt = 4/3(pi) (3r^2)dr/dt dr/dt = 0.16 substitute. Then integrate to get V(t)

OpenStudy (anonymous):

i don't know how to do that

OpenStudy (ranga):

Are you in pre-calculus?

OpenStudy (anonymous):

yes...

OpenStudy (ranga):

We will have to do this without calculus then. Find r as a function of time t. r starts at 38 inches and steadily increases at 0.16 inch per second. So in t seconds r would have increased by 0.16t when added to the initial 38 inches will give: r(t) = 38 + 0.16t Volume of a sphere = 4/3(pi)r^3. Since radius changes with time, volume will change with time too. V(t) = 4/3(pi)r^3 = 4/3(pi)(38 + 0.16t)^3

OpenStudy (ranga):

Put t = 280 seconds and find V.